Cabinaire Inc. is one of the largest manufacturers of office furniture in the United States. In...

  1. Cabinaire Inc. is one of the largest manufacturers of office furniture in the United States. In Grand Rapids, Michigan, it assembles filing cabinets in an Assembly Department. Assume the following information for the Assembly Department:

    Direct labor per filing cabinet 30 minutes
    Supervisor salaries $123,000 per month
    Depreciation $25,000 per month
    Direct labor rate $18 per hour

    Prepare a flexible budget for 9,000, 11,000, and 14,000 filing cabinets for the month ending February 28 in the Assembly Department similar to Exhibit 5. Round your final answers to the whole dollar, if required.

    Cabinaire Inc.
    Assembly Department Budget
    For the Month Ending February 28 (assumed data)
    Units of production 9,000 11,000 14,000
    Variable cost:
    • Depreciation
    • Direct labor
    • Supervisor salaries
    • Rent
    • Utilities
    $ $ $
    Total variable cost $ $ $
    Fixed cost:
    • Direct labor
    • Electric power
    • Rent
    • Supervisor salaries
    • Wages
    $ $ $
    • Depreciation
    • Direct labor
    • Insurance
    • Utilities
    • Wages
    Total fixed cost $ $ $
    Total department costs $ $ $
Answer #1
Assembly Department Budget
For the Month Ending February 28 (assumed data)
Units of production 9000 11000 14000
Variable cost:
Direct labor 81000 99000 126000
Total variable cost 81000 99000 126000
Fixed cost:
Supervisor salaries 123000 123000 123000
Depreciation 25000 25000 25000
Total fixed cost 148000 148000 148000
Total department costs 229000 247000 274000
Workings:
Direct labor =9000*0.5*18 =11000*0.5*18 =14000*0.5*18